An office supply store in San Diego sells 7 writing tablets and 4 pens for $ 6.40. Also, 2 tablets and 19 pens cost # 5.40. Find the price of each.

From a take home test

Answer by checkley77(12844)

7T+4P=6.40
2T+19P=5.40
MULTIPLY THE FIRST EQUATION BY 2 & MULTIPLY THE SECOND EQUATION BY -7 & THEN ADD THEM.
14T+8P=12.80
-14T-133P=-37.80
----------------------
-125P=-25.00
P=-25.00/-125
P=$.20 IS THE PRICE FOR EACH PEN.
7T+4*.20=6.40
7T+.80=6.40
7T=6.40-.80
7T=5.60
T=5.60/7
T=$.80 FOR EACH TABLET.
PROOF:
2*.80 +19*.20=5.40
1.60+3.80=5.40
5.40=5.40